A compound contains arsenic and oxygen as its only elements. A 1.626 g sample of the compound contains 1.060 g of arsenic. What is the empirical formula for the compound?
Answer #1

As mass = 1.060 g

As moles = mass / molar mass

                = 1.060 / 74.92

                = 0.0142

mass of O =1.626 - 1.060 = 0.566

moles of O = 0.566 / 16 = 0.0354

moles ratio of As and O = 0.0142 : 0.0354

                                     = 1 : 2.5

                                     = 2 : 5

empirical formula = As2O5